A person has much of his savings invested in 15,000 shares of Grass Roots common stock. The stock is currently selling for $12 per share and has been paying a dividend of $.75 per share. Grass Roots has discontinued its dividend but begins to grow at 7% a year. Assuming no transaction costs.

Q. How can this person maintain his income and his position in the firm at the end of the year?
Answer #1

Position when Company was giving a dividend

Value of Investment : $180,000 (15000 shares @ $12 per share)

Dividend Income : $11,250 ( $.75 per share for 15,000 shares)

Return on Investment (assuming value of investment = cost of investment) = 11,250/180,000*100 = 6.25%

Position when Company is not paying dividend

Cost of Investment : $180,000

Value of Investment at the end of the year : $180,000 * 1.07 = $192,600

Return on Investment =7%

To maintain the income and his position in the firm:

He can buy a call option of that share at strike price $12 per share and exercise this option when the share reaches at $12.84 (assuming 7% growth).In this way, he can earn the income of $12,600(this amount is before deducting premium to be paid on call option).After deducting premium , he will be able to maintain his income of $11,250.

Another way is he can do the margin trading on the amount in his demat account and earn the income on appreciation of stock.

Please note by income it is assumed it meant realised income.